OpenStudy (anonymous):

The shorter leg of a 30°-60°-90° triangle is 7.4 meters long. Find the perimeter.

OpenStudy (anonymous):

35.02

OpenStudy (mertsj):



OpenStudy (anonymous):

@Mertsj srsly?

OpenStudy (anonymous):

get a life

OpenStudy (anonymous):

Length 7.4 is the side opposite to 30° so hypotenuse = 2 × 7.4 = 14.8 ans side opposite to 60° = √3/2 × 14.8 = √3 × 7.4 perimeter = 7.4 + 14.8 + 7.4 √3 = 22.2 + 7.4√3 m = 35.02

OpenStudy (anonymous):

happy?

OpenStudy (anonymous):

huh?

OpenStudy (anonymous):

@mertsj was reporting me with an inadequate answer.